The bus began its route with some passengers. At the first stop, 4 people got on and 1 got off the bus. At the second stop, 5 people got on and 3 got off. At the third stop, 6 passengers got on and 5 others got off. At the fourth stop, 7 more got on and 3 got off. At the fifth stop, 8 got on and 6 got off. When the bus finally arrived at the terminal, there were 21 passengers on board. How many people were on the bus when it began its route?
Answer by roxy9229(32)

x+4-1+5-3+6-5+7-3+8-6 = 21
x+3+5-3+6-5+7-3+8-6 = 21
x+8-3+6-5+7-3+8-6 = 21
x+5+6-5+7-3+8-6 = 21
x+11-5+7-3+8-6 = 21
x+6+7-3+8-6 = 21
x+13-3+8-6 = 21
x+10+8-6 = 21
x+18-6 = 21
x+12 = 21
-12 -12
x=9
there were 9 people on the bus when it first started